OpenStudy (anonymous):

Find the exact solutions of x^2-(y-6)^2=36 and y=-x^2 ?? I'm getting an imaginary answer, is that correct?

OpenStudy (campbell_st):

well you are looking for the points of intersection between the curves rewrite the 2nd equation as x^2 = -y and substitute into the 1st equation -y - (y -6)^2 = 36 -y -(y^2 - 12y + 36) = 36 so -y^2 + 11y - 72 = 0 divide by -1 gives y^2 -11y + 72 = 0 so yes you get complex solutions, and I think thats why they have asked for the exact solutions hope this helps

OpenStudy (anonymous):

Yes it does! Thanks, I was just looking for a second opinion, I appreciate it!